# UltraClean: A Simple Framework to Train Robust Neural Networks against Backdoor Attacks

### v1

@@ 4.1 Experiment Settings
We consider multiple representative dirty-label and clean-label attacks within the threat model. Please note that model poisoning attacks that disable training phase defense where attackers have full control of training process do not align with our threat model. For dirty-label attacks, we implement BadNets [28], Blended and Random Pattern [10], and Trojan [46] following the original papers, and evaluate the effectiveness of UltraClean on CIFAR-10. For clean-label attacks, we implement Sinusoidal Signal Backdoor (SIG) [2]; Label Consistent Backdoor (LCBD) [62] and Hidden Trigger Backdoor (HTBD) [57] based on their open-source repositories, and adopt the same training algorithms and dataset as in original papers for assessment (i.e., GTSRB for SIG, CIFAR-10 for LCBD, and ImageNet for HTBD). The mechanisms of attacks, detailed training settings, statistics of datasets, and neural network architectures are also summarized in the supplementary materials.
Attack Acc. ASR BDR
Pre-clean (PC)
BadNets 85.27% 100.00% /
Trojan 84.94% 99.19% /
Blended (HK) 84.99% 97.47% /
Blended (RP) 86.24% 99.96% /
UltraClean (UC)
BadNets 83.91% 0.83% 94.42%
Trojan 84.73% 1.61% 99.50%
Blended (HK) 85.08% 3.06% 97.38%
Blended (RP) 84.23% 0.15% 99.80%
Table 1: Performance of UltraClean against dirty-label attacks
Class (ID)
Acc. (PC)
ASR (PC)
Acc. (SVD)
ASR (SVD)
BDR (SVD)
Acc. (UC)
ASR (UC)
BDR (UC)
95.83%
47.58%
96.18%
58.36%
43.48%
96.52%
8.18% 52.17%
97.22%
54.15%
96.18%
62.82%
44.87%
95.49%
10.47% 56.41%
95.83%
67.62%
96.52%
26.33% 67.21%
96.87%
61.92%
50.82%
95.14%
17.38%
97.22%
2.13%
50.99%
95.14%
0.00% 69.09%
95.49%
51.53%
95.49%
40.46%
57.43%
96.18%
1.91%
46.62%
97.57%
51.78%
95.83%
53.75%
42.56%
93.05%
37.15% 52.82%
95.49%
74.16%
96.18%
62.92%
40.16%
96.52%
17.98% 62.99%
95.83%
63.83%
96.88%
62.92%
30.77%
97.22%
55.32% 51.92%
97.92
73.41%
96.18%
75.79%
40.70%
96.18%
40.08% 48.74%
97.57%
47.18%
93.06%
22.54%
64.58%
97.22%
6.34% 91.66%
@@ 4.2 Evaluation on Dirty-Label Attacks
We first demonstrate UltraClean’s effectiveness against dirty-label attacks. We randomly inject 10% poisoned images in the CIFAR-10 dataset and set the target dirty-label to “automobile”. We set the blended injection ratio to 0.2 and Trojan transparency to 0.5, respectively. The results are presented in Table 1. UltraClean achieves an average of 97.78% detection rate and an average of 97.93% ASR reduction at a removal threshold of 0.3, indicating a strong capability of defending against dirty-label attacks. In practice, users usually do not know how many poisoned samples are injected. The selection of $\beta$ should depend on the specific requirement of the application. Under safety-critical scenarios, users can start with a relatively large threshold (e.g., 0.3) and gradually adjust it.
@@ 4.3 Evaluation on Clean-Label Attacks
We are more interested in the performance of UltraClean on clean-label attacks since there is no known effective poisons-filtering-based defense against such stealthy attacks. Thus, we comprehensively evaluate the performance of UltraClean against clean-label attacks. We adapt SVD [61] and STRIP [24] as the baseline methods. For fair comparison, we follow the exactly same settings of SVD and STRIP, and consider two different scenarios:
@@
    detection on the poisoned class (same setting as SVD) where we assume that the defender already knows the poisoned class (i.e., target class);
@@
    detection on the whole training dataset (same setting as STRIP) where the defender does not have knowledge of the data poisoning process.
@@
4.3.1 Detection on the Poisoned Class
@@
HTBD on ImageNet. For the HTBD attack, we follow the same settings in the original paper [57] and conduct the experiment on 10 randomly selected target classes. A total of 100 poisoned samples are injected into the target class. Table 4 illustrates the detection rate, ASR and model accuracy. It can be seen that for most classes, poisoned samples completely bypass the SVD defense. However, UltraClean captures almost all poisoned samples and significantly undermines the backdoor effect to nearly 0% ASR. Moreover, UltraClean even improves the post-clean model accuracy.
@@
Table 4: Comparison of clean accuracy, ASR and BDR against HTBD attack on ImageNet (“*” denotes results replicated from [57]).
@@
Table 5: UltraClean performance on entire dataset
4.3.2 Detection on the Whole Training Dataset
HTBD on ImageNet. The entire ImageNet dataset has more than one million training images. Precisely detecting poisoned samples from such a large-scale dataset is an extremely challenging task. However, UltraClean still acquires considerable success when facing such a complex dataset. In the experiment, we reproduce the attack in the original paper by injecting only 400 poisoned samples ($\sim$0.04of total training samples) into the target “French bulldog” class. The results of UltraClean and STRIP are presented in Table 5 and Figure 2 (right). It can be observed that UltraClean is particularly effective against HTBD, while STRIP is unsuccessful in detecting the poisoned samples. With only a 5% removal threshold, UltraClean achieves a nearly 90% detection rate while maintaining a similar level of model accuracy.
@@ 4.4 Performance on Clean Datasets
Threshold ($\boldsymbol{\beta}$ = 0.30) Regular UltraClean
@@ 4.5 Direct Training on Denoised Datasets
Table 7: Comparison of accuracy and ASR between training on regular poisoned dataset and denoised poisoned dataset (BadNets)
@@ Appendix A Mechanisms of Different Poisons
As discussed in Section 2 in the main paper, various backdoor attacks have been developed in the literature. In this paper, we examine several of the most representative attacks: BadNets [28], blended injection attack [10], Trojan [46], sinusoidal signal backdoor [2], label-consistent backdoor [62], and hidden trigger backdoor [57]. We elucidate the mechanisms of these attacks in this section.
@@ A.1 Dirty-Label Attacks
A.1.1 BadNets
BadNets is the first backdoor attack against deep neural networks, which uses a pattern (e.g., flower, colored square, boom, etc.) as the backdoor trigger. Poisoned samples are patched with the backdoor trigger and assigned an incorrect target label. The poisons generation process are expressed in Equation 7, where $\mathcal{T}$ is the backdoor trigger and $\mathcal{M}$ is the masking area to put the trigger.
@@ A.1.1 BadNets
A.1.2 Blended Injection Attack
Blended crafts backdoor trigger using a fixed cartoon pattern (e.g., Hello Kitty) or a random pattern (e.g., random noise). Similar to BadNets, a target dirty-label is assigned to all poisoned samples. Equation 8 depicts the generation process of Blended, where $\alpha$ is the blended ratio.
@@ A.1.2 Blended Injection Attack
A.1.3 Trojan
Trojan generates the backdoor trigger by assigning value to the input variables in a given trigger mask. The value assignment process aims to make some selected internal neurons (i.e., a neuron in the last layer denoting the target dirty-label) to achieve the maximum values. The neurons selection process can be mathematically expressed by Equation 9 where $L_{target}$ is the layer of target neurons, $L_{preceding}$ is the preceding layers, $W$ is the weights and $b$ is the biases between the layers.
@@ A.1.3 Trojan
Once the dirty-label and corresponding target neurons are selected, the poison samples are generated as shown in Equation 10, where $lr$ is the learning rate and $cost$ is the sum of the square error of target values and actual neuron values.
@@ A.2 Clean-Label Attacks
A.2.1 Sinusoidal Signal Backdoor (SIG)
SIG devises the sinusoidal signal as the backdoor trigger, expressed as:
@@ A.2.1 Sinusoidal Signal Backdoor (SIG)
A.2.2 Label Consistent Backdoor (LCBD)
LCBD places the trigger on images that are hard to classify so that the model relies more on the backdoor trigger to learn. The work proposes two approaches to make images hard-to-classify. The first generative adversarial network (GAN) based method crafts images $\tilde{x_{p}}$ by interpolating a given image $x_{1}$ to another $x_{2}$ through the latent space, where $x_{1}$ is the image from the poisoned class while $x_{2}$ is an arbitrary image from a different class. The process can be expressed as:
$\tilde{z_{1}}=\mathop{\arg\min}\limits_{z_{1}\in\mathds{R}^{d}}||x_{1}-\mathcal{G}(z_{1})||_{2},$ (12)
$\tilde{z_{2}}=\mathop{\arg\min}\limits_{z_{2}\in\mathds{R}^{d}}||x_{2}-\mathcal{G}(z_{2})||_{2},$
$\tilde{x_{p}}=\mathcal{G}({\uptau}\tilde{z_{1}}+(1-{\uptau})\tilde{z_{2}}),$ (13)
@@ A.2.2 Label Consistent Backdoor (LCBD)
The second adversarial perturbation based method leverages the principle of adversarial examples (AE) [27] and constructs hard-to-classify images $\tilde{x_{p}}$ as:
$\tilde{x_{p}}=\mathop{\arg\max}\mathcal{L}(x_{b}+\delta,y,\theta)\quad\text{s.t.}\quad||\delta||_{p}\leqslant\epsilon,$ (14)
@@
A.2.3 Hidden Trigger Backdoor (HTBD)
HTBD is an even stealthier attack based on feature collision. The attack has two steps to generate poisoned data. As shown in Equations (16) and (17), it generates a base image (from the source class) attached with a pre-defined trigger and then optimize poisoned data $x_{p}$ (from the target class) by minimizing their $\ell_{2}$-norm distance with $\tilde{x_{b}}$ in the feature space.
@@ A.2.3 Hidden Trigger Backdoor (HTBD)
$x_{p}=\mathop{\arg\max}\limits_{||x-x_{t}||_{\infty}<\epsilon}||f(x)-f(\tilde{x_{b}})||^{2}_{2}.$ (17)
@@
Appendix B Details of Experimental Settings
We introduce the details of experimental settings in this section. Following the open-sourced repositories of original papers [28, 46, 10, 2, 62, 57], all experiments are implemented in Tensorflow [1] (LCBD [^1] https://github.com/MadryLab/label-consistent-backdoor-code) and PyTorch [52] (BadeNets [^2] https://github.com/Kooscii/BadNets, Blended, Trojan [^3] https://github.com/PurduePAML/TrojanNN, SIG [^4] https://github.com/ebagdasa/backdoor_federated_learning and HTBD [^5] https://github.com/UMBCvision/Hidden-Trigger-Backdoor-Attacks), and run on NVIDIA Tesla V100 GPUs. To reproduce the attacks and achieve the best attack performance, we follow original settings introduced in original papers. The statistics of datasets and models architecture for BadNets, Blended, Trojan, SIG, LCBD, and HTBD are summarized in Table 9. Note that all the datasets used in this paper, i.e., GTSRB [^6] https://benchmark.ini.rub.de/gtsrb_news.html [59], CIFAR-10777https://www.cs.toronto.edu/ kriz/cifar.html, and ImageNet-ILSVRC2012888https://www.image-net.org/challenges/LSVRC/2012/index.php [13], are public.
HTBD. Different from SIG and LCBD, the HTBD trains models in a fine-tuning fashion according to the original paper [57]. During the generation of the poisoned samples, we use the fc7 features of AlexNet (i.e., extract features from the 7th fully-connected layer of AlexNet) for feature collision. Poisoned data are generated with a learning rate of 0.01 in total 2 optimization epochs, and each epoch has 5000 iterations. During fine-tuning, models are trained using SGD with momentum of 0.9 and a learning rate of 0.001. The batch size and epoch are set to 256 and 30 for detection on the poisoned class and 1024 and 10 for the detection on the whole training dataset, respectively.
@@ C.1 Dirty-Label Attacks
We present the ablation experiments results of dirty-label attacks in Table 10. We find that BadNets is more sensitive to the type of baseline images. UltraClean fails to detect poisoned samples only based on baseline images generated by the mean denoising function, while successfully detecting most poisoned samples using baseline images generated by the median denoising function. On the other hand, UltraClean demonstrates consistently superior performance in detecting Blended and Trojan poisoned samples regardless of the type of baseline images. In general, the detection performance reaches the peak when we employ both baseline images. Meanwhile, all three attacks are not very sensitive to the change of the removal threshold. Although higher removal thresholds indeed improve the detection rate, a lower removal threshold can already effectively detect 83$\sim$97poisoned samples.

### v2

@@ 4 Experiments
Class (ID) Acc. (PC) ASR (PC) Acc. (SVD) ASR (SVD) BDR (SVD) Acc. (UC) ASR (UC) BDR (UC)
[Uncaptioned image] (0) 95.83% 47.58% 96.18% 58.36% 43.48% 96.52% 8.18% 52.17%
[Uncaptioned image] (1) 97.22% 54.15% 96.18% 62.82% 44.87% 95.49% 10.47% 56.41%
[Uncaptioned image] (2) 95.83% 67.62% 96.52% 26.33% 67.21% 96.87% 61.92% 50.82%
[Uncaptioned image] (3) 95.14% 17.38% 97.22% 2.13% 50.99% 95.14% 0.00% 69.09%
[Uncaptioned image] (4) 95.49% 51.53% 95.49% 40.46% 57.43% 96.18% 1.91% 46.62%
[Uncaptioned image] (5) 97.57% 51.78% 95.83% 53.75% 42.56% 93.05% 37.15% 52.82%
[Uncaptioned image] (6) 95.49% 74.16% 96.18% 62.92% 40.16% 96.52% 17.98% 62.99%
[Uncaptioned image] (7) 95.83% 63.83% 96.88% 62.92% 30.77% 97.22% 55.32% 51.92%
[Uncaptioned image] (8) 97.92 73.41% 96.18% 75.79% 40.70% 96.18% 40.08% 48.74%
[Uncaptioned image] (9) 97.57% 47.18% 93.06% 22.54% 64.58% 97.22% 6.34% 91.66%
@@ 4.1 Experiment Settings
We consider multiple representative dirty-label and clean-label attacks within the threat model. Please note that attacks that disable training phase defense where attackers train the models do not align with our threat model. For dirty-label attacks, we implement BadNets [gu2017badnets], Blended and Random Pattern [DBLP:journals/corr/abs-1712-05526], and Trojan [Trojannn] following the original papers, and evaluate the effectiveness of UltraClean on CIFAR-10. For clean-label attacks, we implement Sinusoidal Signal Backdoor (SIG) [DBLP:conf/icip/BarniKT19]; Label Consistent Backdoor (LCBD) [DBLP:journals/corr/abs-1912-02771] and Hidden Trigger Backdoor (HTBD) [DBLP:conf/aaai/SahaSP20] based on their open-source repositories, and adopt the same training algorithms and dataset as in original papers for assessment (i.e., GTSRB for SIG, CIFAR-10 for LCBD, and ImageNet for HTBD). The mechanisms of attacks, detailed training settings, statistics of datasets, and neural network architectures are also summarized in the supplementary materials.
@@ 4.2 Evaluation on Dirty-Label Attacks
We first demonstrate UltraClean’s effectiveness against various dirty-label attacks and refer to the SOTA defense against such attacks – the Frequency Detection (FD) [DBLP:conf/iccv/ZengPMJ21] as the baseline for comparison. Note that FD is ineffective against clean-label attacks, as shown in [DBLP:conf/iccv/ZengPMJ21]. We follow the original poisoning practice and set the blended injection ratio to 0.2 and Trojan transparency to 0.5, respectively. The results are presented in Table 2. UltraClean has an average of 97.78% detection rate and an average of 97.93% ASR reduction at a removal threshold of 0.3, achieving comparable performance to FD and indicating a strong capability of defending against dirty-label attacks. In practice, users usually do not know how many poisoned samples are injected. The selection of $\beta$ should depend on the specific requirement of the application. Under safety-critical scenarios, we consider grid search a practical and systematic algorithm to determine the best value of $\beta$ given an accuracy threshold.
Attack
Type
Acc.
ASR
Acc.
ASR
BDR
BDR
BadNets 85.27% 100.00% 83.91% 0.83% 94.42% 90.50%
Trojan 84.94% 99.19% 84.73% 1.61% 99.50% 99.99%
Blended (HK) 84.99% 97.47% 85.08% 3.06% 97.38% 96.30%
Blended (RP) 86.24% 99.96% 84.23% 0.15% 99.80% 96.30%
Table 2: Performance comparison of UltraClean and prior SOTA Frequency Detection against dirty-label attacks.
@@ 4.3 Evaluation on Clean-Label Attacks
We are more interested in the performance of UltraClean on clean-label attacks since there is no known effective poisons-filtering-based defense against such stealthy attacks. Thus, we comprehensively evaluate the performance of UltraClean against clean-label attacks. We adapt SVD [DBLP:conf/nips/Tran0M18] and STRIP [DBLP:conf/acsac/GaoXW0RN19] as the baseline methods for comparison. We follow the works of SVD and STRIP, and consider two different scenarios:
@@
    detection on the poisoned class where we assume that the defender already knows the poisoned class (i.e., target class);
@@
    detection on the whole training dataset where the defender does not have knowledge of the data poisoning process.
@@
4.4 Detection on the Poisoned Class
@@
HTBD on ImageNet. For the HTBD attack, we follow the same settings in the original paper [DBLP:conf/aaai/SahaSP20] and conduct the experiment on 10 randomly selected target classes. A total of 100 poisoned samples are injected into the target class. Table 4 illustrates the detection rate, ASR and model accuracy. It can be seen that for most classes, poisoned samples completely bypass the SVD defense. However, UltraClean captures almost all poisoned samples and significantly undermines the backdoor effect to nearly 0% ASR. Moreover, UltraClean even improves the post-clean model accuracy.
@@
Table 4: Comparison of clean accuracy, ASR and BDR against HTBD attack on ImageNet (“*” denotes results replicated from [DBLP:conf/aaai/SahaSP20]).
4.5 Detection on the Whole Training Dataset
HTBD on ImageNet. The entire ImageNet dataset has more than one million training images. Precisely detecting poisoned samples from such a large-scale dataset is an extremely challenging task. However, UltraClean still acquires considerable success when facing such a complex dataset. In the experiment, we reproduce the attack in the original paper by injecting only 400 poisoned samples ($\sim$0.04% of total training samples) into the target “French bulldog” class. The results of UltraClean and STRIP are presented in Table 5 and Figure 3 (right). It can be observed that UltraClean is particularly effective against HTBD, while STRIP is unsuccessful in detecting the poisoned samples. With only a 5% removal threshold, UltraClean achieves a nearly 90% detection rate while maintaining a similar level of model accuracy.
@@ 4.5 Detection on the Whole Training Dataset
Table 5: UltraClean performance on entire dataset.
@@ 4.6 Performance on Clean Datasets
Threshold
($\boldsymbol{\beta}$ = 0.30)
Regular UltraClean
@@ 4.7 Direct Training on Denoised Datasets
Table 7: Comparison of accuracy and ASR between training on regular poisoned dataset and denoised poisoned dataset (BadNets).
@@ 5 Conclusion
ACKNOWLEDGMENT
This work is supported in part by the National Science Foundation (NSF) under grant numbers 2426299, 2413046, and 2532588.
@@ Appendix A Mechanisms of Different Poisons
As discussed in Section 2 in the main paper, various backdoor attacks have been developed in the literature. In this paper, we examine several of the most representative attacks: BadNets [gu2017badnets], blended injection attack [DBLP:journals/corr/abs-1712-05526], Trojan [Trojannn], sinusoidal signal backdoor [DBLP:conf/icip/BarniKT19], label-consistent backdoor [DBLP:journals/corr/abs-1912-02771], and hidden trigger backdoor [DBLP:conf/aaai/SahaSP20]. We elucidate the mechanisms of these attacks in this section.
@@ A.1 Dirty-Label Attacks
BadNets. BadNets is the first backdoor attack against deep neural networks, which uses a pattern (e.g., flower, colored square, boom, etc.) as the backdoor trigger. Poisoned samples are patched with the backdoor trigger and assigned an incorrect target label. The poisons generation process are expressed in Equation 7, where $\mathcal{T}$ is the backdoor trigger and $\mathcal{M}$ is the masking area to put the trigger.
@@
Blended Injection Attack. Blended crafts backdoor trigger using a fixed cartoon pattern (e.g., Hello Kitty) or a random pattern (e.g., random noise). Similar to BadNets, a target dirty-label is assigned to all poisoned samples. Equation 8 depicts the generation process of Blended, where $\alpha$ is the blended ratio.
@@
Trojan. Trojan generates the backdoor trigger by assigning value to the input variables in a given trigger mask. The value assignment process aims to make some selected internal neurons (i.e., a neuron in the last layer denoting the target dirty-label) to achieve the maximum values. The neurons selection process can be mathematically expressed by Equation 9 where $L_{target}$ is the layer of target neurons, $L_{preceding}$ is the preceding layers, $W$ is the weights and $b$ is the biases between the layers.
@@
Once the dirty-label and corresponding target neurons are selected, the poison samples are generated as shown in Equation 10, where $lr$ is the learning rate and costcost is the sum of the square error of target values and actual neuron values.
@@ A.2 Clean-Label Attacks
Sinusoidal Signal Backdoor (SIG). SIG devises the sinusoidal signal as the backdoor trigger, expressed as:
@@
Label Consistent Backdoor (LCBD). LCBD places the trigger on images that are hard to classify so that the model relies more on the backdoor trigger to learn. The work proposes two approaches to make images hard-to-classify. The first generative adversarial network (GAN) based method crafts images $\tilde{x_{p}}$ by interpolating a given image $x_{1}$ to another $x_{2}$ through the latent space, where $x_{1}$ is the image from the poisoned class while $x_{2}$ is an arbitrary image from a different class. The process can be expressed as:
[…]
@@
The second adversarial perturbation based method leverages the principle of adversarial examples (AE) [DBLP:journals/corr/GoodfellowSS14] and constructs hard-to-classify images $\tilde{x_{p}}$ as:
[…]
@@
Hidden Trigger Backdoor (HTBD). HTBD is an even stealthier attack based on feature collision. The attack has two steps to generate poisoned data. As shown in Equations (16) and (17), it generates a base image (from the source class) attached with a pre-defined trigger and then optimize poisoned data $x_{p}$ (from the target class) by minimizing their $\ell_{2}$-norm distance with $\tilde{x_{b}}$ in the feature space.
@@
[…]
Appendix B Details of Experimental Settings
We introduce the details of experimental settings in this section. Following the open-sourced repositories of original papers [gu2017badnets, Trojannn, DBLP:journals/corr/abs-1712-05526, DBLP:conf/icip/BarniKT19, DBLP:journals/corr/abs-1912-02771, DBLP:conf/aaai/SahaSP20], all experiments are implemented in Tensorflow [tensorflow2015-whitepaper] (LCBD [^1] https://github.com/MadryLab/label-consistent-backdoor-code) and PyTorch [paszke2019pytorch] (BadeNets [^2] https://github.com/Kooscii/BadNets, Blended, Trojan [^3] https://github.com/PurduePAML/TrojanNN, SIG [^4] https://github.com/ebagdasa/backdoor_federated_learning and HTBD [^5] https://github.com/UMBCvision/Hidden-Trigger-Backdoor-Attacks), and run on NVIDIA Tesla V100 GPUs. To reproduce the attacks and achieve the best attack performance, we follow original settings introduced in original papers. The statistics of datasets and models architecture for BadNets, Blended, Trojan, SIG, LCBD, and HTBD are summarized in Table 9. Note that all the datasets used in this paper, i.e., GTSRB [^6] https://benchmark.ini.rub.de/gtsrb_news.html [stallkamp2011german], CIFAR-10777https://www.cs.toronto.edu/ kriz/cifar.html, and ImageNet-ILSVRC2012888https://www.image-net.org/challenges/LSVRC/2012/index.php [deng2009imagenet], are public.
HTBD. Different from SIG and LCBD, the HTBD trains models in a fine-tuning fashion according to the original paper [DBLP:conf/aaai/SahaSP20]. During the generation of the poisoned samples, we use the fc7 features of AlexNet (i.e., extract features from the 7th fully-connected layer of AlexNet) for feature collision. Poisoned data are generated with a learning rate of 0.01 in total 2 optimization epochs, and each epoch has 5000 iterations. During fine-tuning, models are trained using SGD with momentum of 0.9 and a learning rate of 0.001. The batch size and epoch are set to 256 and 30 for detection on the poisoned class and 1024 and 10 for the detection on the whole training dataset, respectively.
@@ C.1 Dirty-Label Attacks
[…]